Supreme01:

The top side of the rectangle is the same as the bottom side of the rectangle. Write an equation and solve to find the value of x.

Supreme01:

Supreme01:

CAN SOMEONE PLEASE HELP ME :)

Gdeinward:

the first step here is to set the equations equal to each other

Gdeinward:

so 3x+18=24+5x. Can you solve this?

Supreme01:

yes i start with the smaller variable and get rid of it correct

Gdeinward:

correct

Supreme01:

ok so i got 18=24+2x

Gdeinward:

correct.

Supreme01:

ok so now do i cover up the variable and the other sign of the equal sign

Gdeinward:

you need to subtract 24 from both sides to isolate the variable

Gdeinward:

so you get -6=2x

Supreme01:

yes i did

Gdeinward:

okay good job. Now all you have to do is divide both sides by 2

Supreme01:

so now i divide -6 by 2

Gdeinward:

yup

Supreme01:

ok so i got -3

Gdeinward:

good job

Supreme01:

ok so i have 5 questions that i did but can you check them

Gdeinward:

yes just make a separate post

Supreme01:

ok thank you

Gdeinward:

oc